Find the volume of the solid that results when the region enclosed by
1
Expert's answer
2020-06-14T17:29:18-0400

Let solid is obtained by "y=\sqrt x" , x-axis and revolve it about "x=9"


then Volume, "V = \int_0^9 \pi r^2 dx = \int_0^9 \pi (\sqrt x)^2 dx = \int_0^9 \pi x dx"


"V = [\frac{\pi}{2}(x^2) ]_0^9 = \frac{\pi}{2}(81-0) = \frac{81 \pi}{2}" unit volume.





Region is bounded by red, blue and green curve is required region.